### v1

# Anticanonical minimal models and Zariski decomposition

Abstract.
Birkar and Hu showed that if a pair $(X,\Delta)$ is lc and $K_{X}+\Delta$ admits a birational Zariski decomposition, then $(X,\Delta)$ has a minimal model. Analogously, we prove that if a pair $(X,\Delta)$ is pklt and $-(K_{X}+\Delta)$ admits a birational Zariski decomposition, then $(X,\Delta)$ has an anticanonical minimal model.
1. Introduction
Ambro ([Amb]) and Fujino ([Fuj11a]) proved the cone theorem and contraction theorem for lc pairs. As a consequence, we can initiate a minimal model program (abbreviated as MMP) for lc pairs. Birkar ([Bir12]) and Hacan, Xu ([HX]) proved the existence of flips for lc pairs. It allows us to keep running the MMP (cf. Remark 3.8). However, we do not know whether the MMP terminates in finite steps. Nevertheless, there are results which support the existence of minimal models. In [BCHM], the authors showed the existence of minimal models for klt pairs of general type. After [BCHM], Birkar proved the existence of minimal models under various conditions (cf. [Bir10, Bir11, Bir12, BH]).
Let $D$ be an $\mathbb{R}$-Cartier divisor on a normal projective variety $X$. We say that $D$ admits a birational Zariski decomposition if there exists a projective birational morphism $f:Y\to X$ such that the positive part $P_{\sigma}(f^{\ast}D)$ of the divisorial Zariski decomposition of $f^{\ast}D$ is nef. It is easy to see that if an lc pair $(X,\Delta)$ has a minimal model, then $K_{X}+\Delta$ admits a birational Zariski decomposition (cf. Section 4.2). In [BH], Birkar and Hu conversely showed that if $K_{X}+\Delta$ admits a birational Zariski decomposition, then $(X,\Delta)$ has a minimal model.
Theorem 1.1 ([BH]).
Let $(X,\Delta)$ an lc pair with $K_{X}+\Delta$ pseudoeffective. If $K_{X}+\Delta$ admits a birational Zariski decomposition, then $(X,\Delta)$ has a minimal model.
In this paper, we are interested in whether the similar result holds when we change the sign of $K_{X}+\Delta$, that is, we replace $K_{X}+\Delta$ with $-(K_{X}+\Delta)$. It is not surprising that this small modification causes the whole argument in the proof to the break down. First of all, it is not at all trivial whether the MMP on $-(K_{X}+\Delta)$ works or not. Furthermore, even if we can run such MMP, the singularities of the resulting model can be arbitrary bad. Thus, we need to control singularities carefully.
Let $(X,\Delta)$ be a pair with $-(K_{X}+\Delta)$ pseudoeffective and let $E$ be a prime divisor over $X$. We define the potential log discrepancy $\bar{a}(E;X,\Delta)$ of $(X,\Delta)$ at $E$ by $\bar{a}(E;X,\Delta):=a(E;X,\Delta)-\sigma_{E}(-(K_{X}+\Delta))$, where $a(E;X,\Delta)$ is the usual log discrepancy and $\sigma_{E}(-(K_{X}+\Delta))$ is the asymptotic valuation. We say that $(X,\Delta)$ is pklt if $\inf_{E}\bar{a}(E;X,\Delta)>0$, where the infimum is taken over all prime divisors $E$ over $X$. These notions were first introduced by Choi and Park in [CP].
Definition 1.2.
Let $(X,\Delta)$ be a pair with $-(K_{X}+\Delta)$ pseudoeffective. We say that a pair $(X^{\prime},\Delta^{\prime})$ is a $-(K_{X}+\Delta)$-minimal model if
  1. $\bullet$
    there exists a birational contraction $\phi:X\dashrightarrow X^{\prime}$, i.e., $\phi$ is a birational map and $\phi^{-1}$ does not contract any divisor,
  2. $\bullet$
    $\Delta^{\prime}=\phi_{\ast}\Delta$ is the birational transform of $\Delta$ on $X^{\prime}$,
  3. $\bullet$
    $(X^{\prime},\Delta^{\prime})$ is $\mathbb{Q}$-factorial dlt,
  4. $\bullet$
    $-(K_{X^{\prime}}+\Delta^{\prime})$ is nef,
  5. $\bullet$
    $\bar{a}(D;X,\Delta)=\bar{a}(D;X^{\prime},\Delta^{\prime})$ for any $\phi$-exceptional prime divisor $D$.
Now, we are ready to introduce our main result.
Theorem 1.3.
Let $(X,\Delta)$ be a pair with $-(K_{X}+\Delta)$ pseudoeffective . Assume that $(X,\Delta)$ is pklt and $-(K_{X}+\Delta)$ admits a birational Zariski decomposition. Then there exists a $-(K_{X}+\Delta)$-minimal model.
Acknowledgement
We thank Professor Sung Rak Choi for helpful suggestions and careful comments on the earlier version of this paper. We also thank Dr. Chuyu Zhou for kind suggestion and corrections.
2. Preliminaries
2.1. Numerical equivalence
Let $X$ be a complete variety. Denote by $\operatorname{Div}_{\mathbb{R}}(X)$ the real vector space generated Cartier divisors on $X$. An element of $\operatorname{Div}_{\mathbb{R}}(X)$ is called $\mathbb{R}$-Cartier divisor and we say that two $\mathbb{R}$-Cartier divisors $D_{1},D_{2}\in\operatorname{Div}_{\mathbb{R}}(X)$ are numerically equivalent, denoted by $D_{1}\equiv D_{2}$, if $D_{1}\cdot C=D_{2}\cdot C$ for every integral curve $C$ on $X$. Let $Z_{1}(X)$ be the real vector space generated by irreducible curves on $X$. An element of $Z_{1}(X)$ is called 1-cycle and we say that two 1-cycles $\gamma_{1},\gamma_{2}\in Z_{1}(X)$ are numerically equivalent, denoted by $\gamma_{1}\equiv\gamma_{2}$, if $D\cdot\gamma_{1}=D\cdot\gamma_{2}$ for every Cartier divisor $D$ on $X$. Define two quotient spaces $N^{1}(X),N_{1}(X)$ by $N^{1}(X):=\operatorname{Div}_{\mathbb{R}}(X)/\equiv$ and $N_{1}(X):=Z_{1}(X)/\equiv$, respectively. Then intersection number gives a perfect pairing
$N^{1}(X)\times N_{1}(X)\to\mathbb{R},\leavevmode\nobreak\ \leavevmode\nobreak\ (\delta,\gamma)\mapsto\delta\cdot\gamma.$
An $\mathbb{R}$-Cartier divisor $D$ on $X$ is nef if $D\cdot C\geq 0$ for any irreducible curve $C$ on $X$. A nef cone $\operatorname{Nef}(X)$ is the cone in $N^{1}(X)$ generated by nef $\mathbb{R}$-Cartier divisors. A Mori cone $\overline{\operatorname{NE}}(X)$ is the cone in $N_{1}(X)$ generated by irreducible curves on $X$. By definition, one can see that $\operatorname{Nef}(X)$ is the dual cone of $\overline{\operatorname{NE}}(X)$.
2.2. Asymptotic valuation
In this section, $X$ denotes a smooth projective variety unless otherwise specified. Let $E$ be a prime divisor on $X$. For a big $\mathbb{R}$-Cartier divisor $D$ on $X$, we define the asymptotic valuation $\sigma_{E}(D)$ of $D$ along $E$ by
[…]
By definition, one can easily see that the following inequality
(2.1) $\sigma_{E}(D_{1}+D_{2})\leq\sigma_{E}(D_{1})+\sigma_{E}(D_{2})$
holds for any big $\mathbb{R}$-divisors $D_{1},D_{2}$ and for any prime divisor $E$ on $X$. Let $A$ be an ample $\mathbb{R}$-divisor on $X$. Then $\sigma_{E}(A)=0$. Indeed, by inequality (2.1), we may assume that $A$ is a very ample divisor on $X$ and obviously $\sigma_{E}(A)=0$ in this case.
Lemma 2.1 ([Nak, Lemma III.1.4]).
Let $E$ be a prime divisor and $D$ a big $\mathbb{R}$-Cartier divisor on $X$. Then
$\sigma_{E}(D)=\lim_{\varepsilon\searrow 0}\sigma_{E}(D+\varepsilon A),$
where $A$ is an ample $\mathbb{R}$-divisor on $X$. In particular, the limit does not depend on the choice of an ample $\mathbb{R}$-divisor.
Proof.
For any $\varepsilon>0$, we have
$\sigma_{E}(D+\varepsilon A)\leq\sigma_{E}(D)+\sigma_{E}(\varepsilon A)=\sigma_{E}(D).$
On the other hand, since $D$ is big, there exist a positive rational number $\delta>0$ and an effective $\mathbb{R}$-divisor $\Delta$ such that $D\sim_{\mathbb{Q}}\delta A+\Delta$ (cf. [Fuj14, Lemma 7.19]). Hence we have
$(1+\varepsilon)\sigma_{E}(D)$ $=\sigma_{E}((1+\varepsilon)D)$
$=\sigma_{E}(D+\varepsilon\delta A+\varepsilon\Delta)$
$\leq\sigma_{E}(D+\varepsilon\delta A)+\varepsilon\operatorname{mult}_{E}\Delta.$
Thus, we obtain that
$\sigma_{E}(D)\leq\liminf\sigma_{E}(D+\varepsilon A)\leq\limsup\sigma_{E}(D+\varepsilon A)\leq\sigma_{E}(D).\qed$
Let $D$ be a pseudoeffective $\mathbb{R}$-divisor and $A^{\prime}$ an ample $\mathbb{R}$-divisor on $X$. Then there exist a positive rational number $\delta$ and an effective $\mathbb{R}$-divisor $\Delta$ such that $A^{\prime}\sim_{\mathbb{Q}}\delta A+\Delta$. Using the similar argument in the proof of Lemma 2.1, we have
$\lim_{\varepsilon\searrow 0}\sigma_{E}(D+\varepsilon A)=\lim_{\varepsilon\searrow 0}\sigma_{E}(D+\varepsilon A^{\prime}).$
Thus the following is well-defined.
Definition 2.2.
Let $E$ be a prime divisor on $X$ and $D$ a pseudoeffective $\mathbb{R}$-Cartier divisor on $X$. Then we define as
$\sigma_{E}(D):=\lim_{\varepsilon\searrow 0}\sigma_{E}(D+\varepsilon A),$
where $A$ is an ample divisor on $X$.
Remark 2.3.
We remark that $\sigma_{E}$ is invariant under numerical equivalence. For a big $\mathbb{R}$-Cartier divisor $D$ on $X$, define $\sigma_{E}(D)_{\text{num}}$ as
$\sigma_{E}(D)_{\text{num}}:=\inf\{\operatorname{mult}_{E}\Delta\leavevmode\nobreak\ |\leavevmode\nobreak\ 0\leq\Delta\equiv D\}.$
Obviously, $0\leq\sigma_{E}(D)_{\text{num}}\leq\sigma_{E}(D)$. Let $A$ be an ample $\mathbb{R}$-divisor and $\Delta$ an effective divisor such that $\Delta\equiv D$. Then $D+\varepsilon A-\Delta$ is an ample $\mathbb{R}$-divisor for any $\varepsilon>0$. Therefore,
$\sigma_{E}(D+\varepsilon A)\leq\operatorname{mult}_{E}\Delta+\sigma_{E}(D+\varepsilon A-\Delta)=\operatorname{mult}_{E}\Delta$
for any $\varepsilon>0$. This shows that $\sigma_{E}(D)_{\text{num}}=\sigma_{E}(D)$. In particular, if $D^{\prime}$ is an $\mathbb{R}$-divisor such that $D\equiv D^{\prime}$, then $\sigma_{E}(D)=\sigma_{E}(D^{\prime})$.
Let $D$ be a pseudoeffective $\mathbb{R}$-Cartier divisor on $X$. The negative part $N_{\sigma}(D)$ of $D$ is defined as
$N_{\sigma}(D)=\sum_{E}\sigma_{E}(D)E$
and the positive part $P_{\sigma}(D)$ of $D$ is defined as $P_{\sigma}(D)=D-N_{\sigma}(D)$. By [Nak, Corollary III.1.11], the above summation is finite. We call $D=P_{\sigma}(D)+N_{\sigma}(D)$ the divisorial Zariski decomposition. If $P_{\sigma}(D)$ is nef, then we call $D=P_{\sigma}(D)+N_{\sigma}(D)$ the Zariski decomposition.
Now, we extend the notion of asymptotic valuation to a singular variety. Let $X$ be a normal projective variety and $D$ a pseudoeffective $\mathbb{R}$-Cartier divisor on $X$. Suppose that $E$ is a prime divisor over $X$. Then there exists a birational morphism $f:Y\to X$ such that $Y$ is a smooth projective variety and $E$ is a prime divisor on $Y$. We define as $\sigma_{E}(D):=\sigma_{E}(f^{\ast}D)=\operatorname{mult}_{E}N_{\sigma}(f^{\ast}D)$. It does not depend on the choice of $f$. Moreover, following [BH], we can define the positive part $P_{\sigma}(D)$ of $D$ as $P_{\sigma}(D)=f_{\ast}P_{\sigma}(f^{\ast}D)$. Similarly, the negative part $N_{\sigma}(D)$ of $D$ is defined as $N_{\sigma}(D)=f_{\ast}N_{\sigma}(f^{\ast}D)$. We also call $D=P_{\sigma}(D)+N_{\sigma}(D)$ the divisorial Zariski decomposition. Note that these definitions do not depend on the choice of $f$. Furthermore, if $P_{\sigma}(D)$ is nef, then the decomposition $D=P_{\sigma}(D)+N_{\sigma}(D)$ is called Zariski decomposition. If in addition $P_{\sigma}(D)$ is semiample, we say that the Zariski decomposition is good. We say that $D$ admits a birational Zariski decomposition if there exists a birational morphism $f:Y\to X$ such that $Y$ is a smooth projective variety and $P_{\sigma}(f^{\ast}D)$ is nef.
2.3. Movable divisors
Let $X$ be a normal projective variety and $D$ a pseudoeffective $\mathbb{R}$-Cartier divisor on $X$. We say that $D$ is movable if $N_{\sigma}(D)=0$. We also say that a class in $N^{1}(X)$ is movable if it is represented by movable divisor. Denote by $\overline{\operatorname{Mov}}(X)$ the closure of the cone generated by movable classes in $N^{1}(X)$. Then by [Nak, Lemma 1.7], $\overline{\operatorname{Mov}}(X)$ coincides with the closure of the convex hull of fixed part free divisors $D$, i.e. $\operatorname{Fix}|D|=0$.
2.4. Potential log discrepancy
Let $X$ be a normal projective variety. An effective $\mathbb{R}$-divisor $\Delta$ on $X$ is said to be a boundary divisor if its coefficients are less than or equal to 1. A pair $(X,\Delta)$ consists of a normal projective variety $X$ and a boundary $\mathbb{R}$-divisor $\Delta$ on $X$ such that $K_{X}+\Delta$ is $\mathbb{R}$-Cartier. If we allow $\Delta$ to have negative coefficients, then $(X,\Delta)$ is called sub-pair.
Let $(X,\Delta)$ be a pair and $E$ a prime divisor over $X$. Then there is a projective birational morphism $f:Y\to X$ such that $Y$ is smooth and $E$ is a divisor on $Y$. The log discrepancy $a(E;X,\Delta)$ of $(X,\Delta)$ at $E$ is defined as $a(E;X,\Delta)=1+\operatorname{mult}_{E}(K_{Y}-f^{\ast}(K_{X}+\Delta))$. We say that $(X,\Delta)$ is klt if $\inf_{E}a(E;X,\Delta)>0$, where the infimum is taken over all prime divisor $E$ over $X$. The log discrepancy is also defined for a sub-pair by the same way and we say that a sub-pair $(X,\Delta)$ is sub-klt if it satisfies the same inequality above.
Let $(X,\Delta)$ be a pair with pseudoeffective $\mathbb{R}$-divisor $-(K_{X}+\Delta)$ and $E$ a prime divisor over $X$. The potential log discrepancy $\bar{a}(E;X,\Delta)$ of $(X,\Delta)$ at $E$ is defined as
$\bar{a}(E;X,\Delta)=a(E;X,\Delta)-\sigma_{E}(-(K_{X}+\Delta)),$
where $\sigma_{E}(-(K_{X}+\Delta))$ is the asymptotic valuation. We say that $(X,\Delta)$ is pklt if $\inf_{E}\bar{a}(E;X,\Delta)>0$, where the infimum is taken over all prime divisor $E$ over $X$. The notion of potential log discrepancy is first introduced and studied in [CP].
Let $\phi:X\dashrightarrow X^{\prime}$ be a birational contraction of normal projective varieties, $D$ an $\mathbb{R}$-Cartier divisor on $X$, and let $D^{\prime}=\phi_{\ast}D$ be the birational transform of $D$. We say that $\phi$ is $D$-nonpositive if there exists a common log resolution $p:W\to X$ and $q:W\to X^{\prime}$ and we can write
$p^{\ast}D=q^{\ast}D^{\prime}+E,$
where $E$ is an effective $q$-exceptional divisor. The potential log discrepancy of $(X,\Delta)$ is preserved along a $-(K_{X}+\Delta)$-nonpositive map. It is a one of significant properties of potential log discrepancy.
Proposition 2.4 ([CP, Proposition 3.11]).
Let $(X,\Delta)$ a pair with pseudoeffective $\mathbb{R}$-divisor $-(K_{X}+\Delta)$. Suppose that there is a $-(K_{X}+\Delta)$-nonpositive map $\phi:X\dashrightarrow X^{\prime}$, and let $\Delta^{\prime}$ be the birational transform of $\Delta$ on $X^{\prime}$. Then we have
$\bar{a}(E;X,\Delta)=\bar{a}(E;X^{\prime},\Delta^{\prime})$
for any prime divisor $E$ over $X$.
3. Negative parts of the divisorial Zariski decompositions
In this section, we will investigate the properties of the negative part of the divisorial Zariski decomposition. To do so, we need a general negativity lemma.
3.1. General negativity lemma
A projective morphism $f:X\to Z$ of quasi-projective varieties is called a contraction if $f_{\ast}\mathcal{O}_{X}=\mathcal{O}_{Z}$. Let $f:X\to Z$ be a contraction. We say that a Cartier divisor $D$ on $X$ is $f$-nef (or nef/$Z$) if $D\cdot C\geq 0$ for every irreducible curve such that $f(C)$ is a point in $Z$.
Lemma 3.1 (cf.[KM, Lemma 3.39]).
Let $f:X\to Z$ be a proper birational morphism of normal varieties. Let $B$ be a $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor on $X$ such that $-B$ is $f$-nef. Then $B$ is effective if and only if $f_{\ast}B$ is effective.
The above lemma, called the negativity lemma, is useful in many situations. There are generalizations of the negativity lemma (cf. [Sho, Lemma 3.22], [Prok, Lemma 1.7]).
Definition 3.2.
Let $f:X\to Z$ be a contraction of normal varieties, $D$ an $\mathbb{R}$-Cartier divisor on $X$, and $V\subseteq X$ a closed subset. We say that $V$ is vertical over $Z$ if $f(V)$ is a proper subset of $Z$. We say that $D$ is very exceptional over $Z$ if $D$ is vertical/$Z$ and for any prime divisor $P$ on $Z$, there exists a prime divisor $Q$ on $X$ such that $Q$ is not a component of $D$ and $f(Q)=P$, i.e., over the generic point of $P$, we have $\operatorname{Supp}f^{\ast}P\not\subseteq\operatorname{Supp}D$.
If $\operatorname{codim}f(D)\geq 2$, then $D$ is very exceptional. On the other hand, when $f$ is birational, then the notions of exceptional and very exceptional coincide.
Let $S\to Z$ be a projective morphism of varieties and $M$ an $\mathbb{R}$-Cartier divisor on $S$. We say that $M$ is nef on the very general curves of $S/Z$ if there is a countable union $\Lambda$ of proper closed subsets of $S$ such that $M\cdot C\geq 0$ for any curve $C$ on $S$ contracted over $Z$ satisfying $C\not\subseteq\Lambda$.
The following is a generalization of the negativity lemma:
Lemma 3.3 ([Sho, Lemma 3.22]).
Let $f:X\to Z$ be a contraction of normal varieties. Let $D$ be an $\mathbb{R}$-Cartier divisor on $X$ written as $D=D^{+}-D^{-}$ with $D^{+},D^{-}\geq 0$ having no common components. Assume that $D^{-}$ is very exceptional/$Z$, and that for each component $S$ of $D^{-}$, $-D|_{S}$ is nef on the very general curves of $S/Z$. Then $D\geq 0$.
For the proof of the above lemma, we refer to [Bir12].
3.2. Behavior of the negative part
Following lemmas describe the behavior of the negative part of a divisorial Zarsiki decomposition. All lemmas have their origins in [Nak] or [BH]. For the convenience, we presents the proofs here.
Lemma 3.4 (cf.[Nak, Theorem III.5.16]).
Let $X$ be a normal projective variety and $D$ a pseudoeffective $\mathbb{R}$-divisor on $X$. Let $f:Y\to X$ be a birational morphism from a smooth projective variety $Y$. Then there exists an effective $f$-exceptional divisor $\Gamma$ on $Y$ such that
$N_{\sigma}(f^{\ast}D)=f^{\ast}N_{\sigma}(D)+\Gamma.$
Furthermore, if $P_{\sigma}(D)$ is nef, then $N_{\sigma}(f^{\ast}D)=f^{\ast}N_{\sigma}(D)$.
Proof.
It is an immediate consequence of Lemma 3.3. Note that $N_{\sigma}(f^{\ast}D)-f^{\ast}N_{\sigma}(D)$ is $f$-exceptional and
$f^{\ast}(D-N_{\sigma}(D))=P_{\sigma}(f^{\ast}D)+N_{\sigma}(f^{\ast}D)-f^{\ast}N_{\sigma}(D).$
By Lemma 3.3, $\Gamma:=N_{\sigma}(f^{\ast}D)-f^{\ast}N_{\sigma}(D)$ is effective.
Assume that $P_{\sigma}(D)$ is nef. Then $f^{\ast}P_{\sigma}(D)$ is also nef and $f^{\ast}N_{\sigma}(D)\geq N_{\sigma}(f^{\ast}D)$ by definition. Hence, $f^{\ast}N_{\sigma}(D)=N_{\sigma}(f^{\ast}D)$. ∎
Lemma 3.5 ([Nak, Lemma III.1.8]).
Let $X$ be a smooth projective variety and $D$ a pseudoeffective $\mathbb{R}$-divisor on $X$. Assume that $0\leq N\leq N_{\sigma}(D)$. Then $N_{\sigma}(D-N)=N_{\sigma}(D)-N$.
Proof.
When $D$ is big, the assertion immediately follows from the definition (cf. [Nak, Lemma III.1.4-(4)]). Now, assume that $D$ is pseudoeffective. Let $A$ be an ample $\mathbb{R}$-divisor on $X$ and let $10<\varepsilon<$ 1. Then for all sufficiently small $\delta>0$,
$N_{\sigma}(D+\delta A)\geq(1-\varepsilon)N_{\sigma}(D).$
Since $D+\delta A$ is big, we have
$N_{\sigma}(D+\delta A-(1-\varepsilon)N)=N_{\sigma}(D+\delta A)-(1-\varepsilon)N.$
Thus,
$N_{\sigma}(D-(1-\varepsilon)N)=N_{\sigma}(D)-(1-\varepsilon)N.$
Letting $\varepsilon\to 0$, by the lower semicontinuity ([Nak, Lemma III.1.7]), we have
$N_{\sigma}(D-N)\leq N_{\sigma}(D)-N.$
On the other hand, by the convexity (2.1),
$N_{\sigma}(D)\leq N_{\sigma}(D-N)+N.$
Hence, we have the desired result. ∎
Remark 3.6.
Let $X$ be a smooth projective variety and $D$ a pseudoeffective $\mathbb{R}$-divisor on $X$. Denote by $P=P_{\sigma}(D),N=N_{\sigma}(D)$. Then for any $x\geq 0$, we have
$N_{\sigma}(D+xP)=N_{\sigma}((1+x)D-xN)=(1+x)N-xN=N.$
Lemma 3.7 ([Nak, Lemma III.5.14]).
Let $X$ be a normal projective variety and $f:Y\to X$ a resolution of $X$. Let $D$ be an $\mathbb{R}$-Cartier divisor on $X$ and $E$ an effective $f$-exceptional divisor. Then
$N_{\sigma}(f^{\ast}D+E)=N_{\sigma}(f^{\ast}D)+E.$
Proof.
Consider the decomposition
$f^{\ast}D=P_{\sigma}(f^{\ast}D+E)+N_{\sigma}(f^{\ast}D+E)-E.$
By the general negativity lemma, $N_{\sigma}(f^{\ast}D+E)-E\geq 0$. Hence, by Lemma 3.5, we can conclude that
$N_{\sigma}(f^{\ast}D)=N_{\sigma}(f^{\ast}D+E)-E.$
Remark 3.8 ($D$-MMP).
Let $X$ be a normal projective variety and $D$ an $\mathbb{R}$-Cartier divisor on $X$. Let $R$ be a $D$-negative extremal ray, i.e., $R$ is an extremal ray of $\overline{\operatorname{NE}}(X)$ such that $D\cdot R<0$. We say that a contraction $f:X\to Z$ is a $D$-negative extremal contraction if it satisfies the following conditions:
  1. $\bullet$
    for any curve $C$ on $X$, $f(C)=\operatorname{pt}$ if and only if $[C]\in R$, and
  2. $\bullet$
    $\rho(X/Z)=1$,
where $\rho(X/Z)$ denotes relative picard number of $X$ over $Z$. Assume that there exists a $D$-negative extremal contraction $f:X\to Z$. Consider the following three possibilities:
  1. (1)
    (fibration) $\dim X>\dim Z$.
  2. (2)
    (divisorial contraction) $f$ is birational and $\operatorname{Exc}(f)$ is a divisor.
  3. (3)
    (flipping contraction) $f$ is birational and $\operatorname{Exc}(f)$ has codimension $\geq 2$.
We proceed the $D$-MMP as follows: If $f:X\to Z$ is a fibration, then we stop. If $f:X\to Z$ is a divisorial contraction, then we let $X^{\prime}=Z$ and $\phi=f$. If $f:X\to Z$ is a flipping contraction, then assuming the existence of the $D$-flip $f^{\prime}:X^{\prime}\to Z$ of $f$ (cf. [Bir07, Definition 2.3]), we have a birational map $\phi:X\dashrightarrow X^{\prime}$. In any case, there is a birational map $\phi:X\dashrightarrow X^{\prime}$. Let $D^{\prime}$ be the birational transform of $D$ on $X^{\prime}$. Assume furthermore that $D^{\prime}$ is $\mathbb{R}$-Cartier (for instance, it is true if $X^{\prime}$ is $\mathbb{Q}$-factorial). If $D^{\prime}$ is nef, then we stop. Otherwise, choose a $D^{\prime}$-negative extremal ray $R^{\prime}$ of $\overline{\operatorname{NE}}(X^{\prime})$ and continue this process on $X^{\prime}$ with $D^{\prime}$. In this way, we obtain a sequence of birational maps
$X=X_{1}\dashrightarrow X_{2}\dashrightarrow\cdots$
which is called a $D$-MMP.
The following lemma shows how positive part of divisorial Zariski decomposition behaves along the MMP.
Lemma 3.9 (cf.[BH, Lemma 4.1]).
Let $X$ be a normal projective variety and $D$ a pseudoeffective $\mathbb{R}$-divisor on $X$. If $\phi:X\dashrightarrow X^{\prime}$ is the $D$-MMP, then $\phi_{\ast}P_{\sigma}(D)=P_{\sigma}(\phi_{\ast}D)$.
Proof.
Let $p:Y\to X$ and $q:Y\to X^{\prime}$ be a common resolution. Write
$p^{\ast}D=q^{\ast}\phi_{\ast}D+E.$
By assumption, $E$ is a $q$-exceptional effective divisor. By Lemma 3.7,
$N_{\sigma}(p^{\ast}D)=N_{\sigma}(q^{\ast}\phi_{\ast}D+E)=N_{\sigma}(q^{\ast}\phi_{\ast}D)+E$
which implies that $P_{\sigma}(p^{\ast}D)=P_{\sigma}(q^{\ast}\phi_{\ast}D)$. ∎
4. Anticanonical minimal models and Zariski decompositions
4.1. MMP for a polarized pair
Let $(X,B)$ be a pair and $P$ a nef $\mathbb{R}$-divisor on $X$. Then $(X,B+P)$ is called a polarized pair. Birkar and Hu first introduced a polarized pair in [BH]. They established the MMP for a polarized pair and proved the existence of minimal model for an lc pair $(X,B)$ with the assumption that $K_{X}+B$ has a birational Zariski decomposition.
A polarized pair $(X,B+P)$ is said to be lc (resp. klt) if $(X,B)$ is lc (resp. klt). Let $(X,B+P)$ be a polarized lc pair and let $R$ be a $(K_{X}+B+P)$-negative extremal ray. For a sufficiently small ample $\mathbb{R}$-divisor $A$ on $X$, there exists a boundary divisor $\Delta$ such that $(X,\Delta)$ is lc, $(K_{X}+\Delta)\cdot R<0$, and
$K_{X}+\Delta\sim_{\mathbb{R}}K_{X}+B+P+A.$
By the contraction theorem for lc pairs, there is a $(K_{X}+\Delta)$-negative extremal contraction $f:X\to Z$. Moreover, if $f$ is a flipping contraction, then there exists a flip by [Bir12, HX].
Suppose that $\phi:X\dashrightarrow X^{\prime}$ is either a divisorial contraction or a flip. Let $K_{X^{\prime}}+B^{\prime}+P^{\prime}+A^{\prime}$ be the birational transform of $K_{X}+B+P+A$ and let $R^{\prime}$ be a $(K_{X^{\prime}}+B^{\prime}+P^{\prime})$-negative extremal ray. Then replacing $A$ if necessary, we can find a divisor $\Delta^{\prime}$ on $X^{\prime}$ such that $(X^{\prime},\Delta^{\prime})$ is lc and
$K_{X^{\prime}}+\Delta^{\prime}\sim_{\mathbb{R}}K_{X^{\prime}}+B^{\prime}+P^{\prime}+A^{\prime}.$
In this way, we can run a $(K_{X}+B+P)$-MMP (cf. [BH, Section 3]).
Remark 4.1.
Let $(X,B+P)$ be a klt polarized pair with $K_{X}+B+P$ pseudoeffective and let $\phi_{i}:X_{i}\dashrightarrow X_{i+1}$ be a $(K_{X}+B+P)$-MMP with scaling of an ample divisor $A$ on $X$, with $X_{1}=X$. After finitely many steps, the MMP consists of only flips. That is, there exists a positive integer $i$ such that $\phi_{j}:X_{j}\dashrightarrow X_{j+1}$ is a flip for all $j\geq i$. Assume that the MMP terminates with a model $X^{\prime}$. Then the birational transform $K_{X^{\prime}}+B^{\prime}+P^{\prime}$ of $K_{X}+B+P$ is nef and
$N_{\sigma}(K_{X^{\prime}}+B^{\prime}+P^{\prime})=0.$
If there are infinitely many flips, then $N_{\sigma}(K_{X_{i}}+B_{i}+P_{i})=0$, where $K_{X_{i}}+B_{i}+P_{i}$ is the birational transform of $K_{X}+B+P$ (cf. [Fuj11b, Theorem 2.3]). Thus, in any case, we arrive at a model $X^{\prime}$ where $N_{\sigma}(K_{X^{\prime}}+B^{\prime}+P^{\prime})=0$.
A $(K_{X}+B+P)$-MMP is said to be $P$-trivial if $P$ is numerically trivial on each extremal ray appear in the MMP.
Theorem 4.2 ([BH, Theorem 3.2]).
Let $(X,B+P)$ be a $\mathbb{Q}$-factorial dlt polarized pair. Assume either that $P$ or $N:=K_{X}+B-P$ is a $\mathbb{Q}$-divisor, or that $N\geq 0$ with $\operatorname{Supp}N\subseteq\operatorname{Supp}B$. Then, for any $\alpha\gg 1$, any $(K_{X}+B+\alpha P)$-MMP is $P$-trivial.
Definition 4.3.
A pair $(Y,B_{Y})$ is a minimal model of an lc pair $(X,B)$ if
  1. $\bullet$
    there exist a birational map $\phi:X\dashrightarrow Y$ and $B_{Y}=\phi_{\ast}B+E$, where $E$ is the sum of the $\phi^{-1}$-exceptional prime divisors on $Y$,
  2. $\bullet$
    $(Y,B_{Y})$ is $\mathbb{Q}$-factorial dlt,
  3. $\bullet$
    $K_{Y}+B_{Y}$ is nef,
  4. $\bullet$
    $a(D;X,B)<a(D;Y,B_{Y})$ for any $\phi$-exceptional prime divisor $D$ on $X$.
Theorem 4.4 ([BH, Theorem 1.1]).
Let $(X,B)$ an lc pair with $K_{X}+B$ pseudoeffective. If $K_{X}+B$ admits a birational Zariski decomposition, then $(X,B)$ has a minimal model.
4.2. Anticanonical minimal models
The converse of Theorem 1.3 is actually easy to prove. Let $(X,\Delta)$ be a pair with $-(K_{X}+\Delta)$ a pseudoeffective $\mathbb{R}$-Cartier divisor. Assume that there exists a $-(K_{X}+\Delta)$-minimal model $(X^{\prime},\Delta^{\prime})$ with a birational map $\phi:X\dashrightarrow X^{\prime}$. Let $p:V\to X$ and $q:V\to X^{\prime}$ be a common log resolution. Write
$-p^{\ast}(K_{X}+\Delta)=-q^{\ast}(K_{X^{\prime}}+\Delta^{\prime})+\Phi.$
Then $\Phi=\sum_{E}(a(E;X,\Delta)-a(E;X^{\prime},\Delta^{\prime}))E$, where $E$ is a prime divisor on $V$. If $E$ is a divisor on $X$, then by definition,
$a(E;X,\Delta)-a(E;X^{\prime},\Delta^{\prime})=a(E;X,\Delta)-\bar{a}(E;X,\Delta)=\sigma_{E}(-(K_{X}+\Delta))\geq 0.$
Therefore, $p_{\ast}\Phi\geq 0$ and by Lemma 3.1, $\Phi\geq 0$. Since $\Phi$ is $q$-exceptional divisor, by Lemma 3.7,
($\ast$) $N_{\sigma}(-p^{\ast}(K_{X}+\Delta))=N_{\sigma}(-q^{\ast}(K_{X^{\prime}}+\Delta^{\prime})+\Phi)=\Phi.$
Hence, we have $P_{\sigma}(-p^{\ast}(K_{X}+\Delta))=-p^{\ast}(K_{X}+\Delta)-\Phi=-q^{\ast}(K_{X^{\prime}}+\Delta^{\prime})$ is nef. Namely, $-(K_{X}+\Delta)$ admits a birational Zariski decomposition.
Remark 4.5.
Let $(X^{\prime},\Delta^{\prime})$ be a $-(K_{X}+\Delta)$-minimal model. Then by the equation ($\ast$4.2), we have
$\bar{a}(E;X,\Delta)=\bar{a}(E;X^{\prime},\Delta^{\prime})$
for any prime divisor $E$ over $X$.
Before starting the proof, we fix some notation for the convenience. Let $\phi:Y\dashrightarrow Y^{\prime}$ be a birational morphism of projective varieties. For an $\mathbb{R}$-divisor $D_{Y}$ on $Y$, we denote by $D_{Y^{\prime}}$ the birational transform of $D_{Y}$ on $Y^{\prime}$ via $\phi$.
Proof of Theorem 1.3.
Let $f:Y\to X$ be a log resolution such that $-f^{\ast}(K_{X}+\Delta)$ admits the Zariski decomposition $-f^{\ast}(K_{X}+\Delta)=P_{Y}+N_{Y}$. Let $K_{Y}+\Delta_{Y}=f^{\ast}(K_{X}+\Delta)$. Then we have
$K_{Y}+\Delta_{Y}+N_{Y}+P_{Y}=0$
and $\bar{a}(E;X,\Delta)=1-\operatorname{mult}_{E}(\Delta_{Y}+N_{Y})$ for a prime divisor $E$ on $Y$. By assumption, the sub-pair $(Y,\Delta_{Y}+N_{Y})$ is sub-klt. Choose a sufficiently small positive real number $\varepsilon>0$ such that $(Y,\Delta_{Y}+(1+\varepsilon)N_{Y})$ is sub-klt.
By taking a higher model if necessary, we may assume that $\operatorname{Exc}(f)\cup\operatorname{Supp}f_{\ast}^{-1}\Delta\cup\operatorname{Supp}N_{Y}$ is a reduced simple normal crossing divisor. Let $E_{1},\dots,E_{k}$ be exceptional divisors of $f$, and write $\Delta_{Y}=f_{\ast}^{-1}\Delta+\sum_{i=1}^{k}d_{i}E_{i}$. For sufficiently small real number $e>0$, define a $f$-exceptional divisor $\Gamma$ as
$\operatorname{mult}_{E_{i}}\Gamma:=\begin{cases}-d_{i}&\leavevmode\nobreak\ \text{if $d_{i}<0$},\\ e&\leavevmode\nobreak\ \text{otherwise}\end{cases}$
and let $B_{Y}$ be a divisor on $Y$ defined as
(4.1) $B_{Y}=\Delta_{Y}+(1+\varepsilon)N_{Y}+\Gamma.$
Denote by $P,N$ the birational transforms of $P_{Y},N_{Y}$, respectively, and let $B=\Delta+(1+\varepsilon)N$. Then one can see that the followings hold:
  1. $\bullet$
    $K_{Y}+B_{Y}+(1+\varepsilon)P_{Y}=f^{\ast}(K_{X}+B+(1+\varepsilon)P)+\Gamma$,
  2. $\bullet$
    $(Y,B_{Y})$ is a klt pair,
  3. $\bullet$
    $\operatorname{Supp}\Gamma=\operatorname{Exc}(f)$.
By Lemma 3.7, we have
$N_{\sigma}(K_{Y}+B_{Y}+(1+\varepsilon)P_{Y})$ $=N_{\sigma}(f^{\ast}(K_{X}+B+(1+\varepsilon)P)+\Gamma)$
$=N_{\sigma}(f^{\ast}(K_{X}+B+(1+\varepsilon)P))+\Gamma$
(4.2) $=\varepsilon N_{Y}+\Gamma.$
Hence the positive part is given as
(4.3) $P_{\sigma}(K_{Y}+B_{Y}+(1+\varepsilon)P_{Y})=\varepsilon P_{Y}.$
Let $A_{Y}$ be an ample divisor on $Y$ and $\alpha>0$ a positive real number. Since $(Y,B_{Y})$ is klt and $P_{Y}$ is nef, we can run a MMP for $K_{Y}+B_{Y}+\alpha P_{Y}$ with scaling of $A_{Y}$. By Theorem 4.2, such MMP is $P_{Y}$-trivial for sufficiently large $\alpha\gg 1$. That is, the nefness of $P_{Y}$ is preserved through the MMP. By Remark 4.1, we arrive at a model $Y^{\prime}$ such that
(4.4) $N_{\sigma}(K_{Y^{\prime}}+B_{Y^{\prime}}+\alpha P_{Y^{\prime}})=0.$
Now, we will trace the positive part of $K_{Y}+B_{Y}+\alpha P_{Y}$. By Remark 3.6, the positive part is given as
$P_{\sigma}(K_{Y}+B_{Y}+\alpha P_{Y})$ $=P_{\sigma}(K_{Y}+B_{Y}+(1+\varepsilon)P_{Y}+(\alpha-1-\varepsilon)P_{Y})$
$=(\alpha-1)P_{Y}.$
Let $\phi:Y\dashrightarrow Y^{\prime}$ be the MMP. By Lemma 3.9, we obtain that
$P_{\sigma}(K_{Y^{\prime}}+B_{Y^{\prime}}+\alpha P_{Y^{\prime}})=\phi_{\ast}P_{\sigma}(K_{Y}+B_{Y}+\alpha P_{Y})=(\alpha-1)P_{Y^{\prime}}.$
On the other hand, by (4.4), we have
$P_{\sigma}(K_{Y^{\prime}}+B_{Y^{\prime}}+\alpha P_{Y^{\prime}})$ $=K_{Y^{\prime}}+B_{Y^{\prime}}+\alpha P_{Y^{\prime}}$
$=(\alpha-1)P_{Y^{\prime}}+\varepsilon N_{Y^{\prime}}+\Gamma_{Y^{\prime}}.$
Therefore, $\varepsilon N_{Y^{\prime}}+\Gamma_{Y^{\prime}}=0$ and $K_{Y^{\prime}}+B_{Y^{\prime}}+P_{Y^{\prime}}=0$.
Let $\psi=\phi\circ f^{-1}:X\dashrightarrow Y^{\prime}$ and let $\Delta_{Y^{\prime}}=\psi_{\ast}\Delta$ be the birational transform of $\Delta$ on $Y^{\prime}$. We will show that $(Y^{\prime},\Delta_{Y^{\prime}})$ is a $-(K_{X}+\Delta)$-minimal model. Since $\operatorname{Supp}\Gamma=\operatorname{Exc}(f)$ and $\Gamma_{Y^{\prime}}=0$, $\psi:X\dashrightarrow Y^{\prime}$ is a birational contraction. Note that
$B_{Y^{\prime}}=\Delta_{Y^{\prime}}+(1+\varepsilon)N_{Y^{\prime}}+\Gamma_{Y^{\prime}}=\Delta_{Y^{\prime}}.$
Thus $-(K_{Y^{\prime}}+\Delta_{Y^{\prime}})=P_{Y^{\prime}}$ is nef. By construction, the $(K_{Y}+B_{Y}+\alpha P_{Y})$-MMP is also $(K_{Y}+B_{Y})$-MMP. It follows that $(Y^{\prime},B_{Y^{\prime}}=\Delta_{Y^{\prime}})$ is $\mathbb{Q}$-factorial klt. We have shown that $(Y^{\prime},\Delta_{Y^{\prime}})$ satisfies first four conditions to be a $-(K_{X}+\Delta)$-minimal model. To complete the proof, we need to compare potential log discrepancies $\bar{a}(E;X,\Delta)$ and $\bar{a}(E;Y^{\prime},\Delta_{Y^{\prime}})$ for a prime divisor $E$ on $X$.
Let $p:V\to Y$ and $q:V\to Y^{\prime}$ be a common log resolution. Then there is an effective $q$-exceptional diviosr $\Phi$ such that
$p^{\ast}(K_{Y}+B_{Y}+\alpha P_{Y})=q^{\ast}(K_{Y^{\prime}}+B_{Y^{\prime}}+\alpha P_{Y^{\prime}})+\Phi.$
By (4.2) and Lemma 3.4, we have
$N_{\sigma}(p^{\ast}(K_{Y}+B_{Y}+\alpha P_{Y}))=p^{\ast}(\varepsilon N_{Y}+\Gamma).$
Since $q^{\ast}(K_{Y^{\prime}}+B_{Y^{\prime}}+\alpha P_{Y^{\prime}})=(\alpha-1)q^{\ast}P_{Y^{\prime}}$ is nef and $\Phi$ is an effective $q$-exceptional divisor, it follows from Lemma 3.7 that
(4.5) $\Phi=p^{\ast}(\varepsilon N_{Y}+\Gamma)$
and
$(\alpha-1)p^{\ast}P_{Y}$ $=P_{\sigma}(p^{\ast}(K_{Y}+B_{Y}+\alpha P_{Y}))$
$=P_{\sigma}(q^{\ast}(K_{Y^{\prime}}+B_{Y^{\prime}}+\alpha P_{Y^{\prime}}))$
$=(\alpha-1)q^{\ast}P_{Y^{\prime}}.$
Therefore $p^{\ast}P_{Y}=q^{\ast}P_{Y^{\prime}}$ and $\Phi=\sum_{E}(a(E;Y^{\prime},B_{Y^{\prime}})-a(E;Y,B_{Y}))E$, where $E$ runs over prime divisors on $V$. Let $E$ be a prime divisor on $Y$. Then we have
$a(E;Y,B_{Y})$ $=1-\operatorname{mult}_{E}B_{Y}$
$=a(E;X,\Delta)-(1+\varepsilon)\operatorname{mult}_{E}N_{Y}-\operatorname{mult}_{E}\Gamma$
$=\bar{a}(E;X,\Delta)-\operatorname{mult}_{E}(\varepsilon N_{Y}+\Gamma).$
On the other hand, by (4.5), we have
$a(E;Y^{\prime},\Delta_{Y^{\prime}})$ $=a(E;Y^{\prime},B_{Y^{\prime}})$
$=a(E;Y,B_{Y})+\operatorname{mult}_{E}(\varepsilon N_{Y}+\Gamma).$
Hence, we can conclude that
$\bar{a}(E;X,\Delta)=a(E;Y,\Delta_{Y^{\prime}})$
for any prime divisor $E$ on $Y$. ∎
Remark 4.6.
Let $(X,\Delta)$ be a pair with pseudoeffective $\mathbb{R}$-divisor $-(K_{X}+\Delta)$. We can show that the existence of $-(K_{X}+\Delta)$-minimal model for movable $\mathbb{R}$-divisor $-(K_{X}+\Delta)$ implies the existence of $-(K_{X}+\Delta)$-minimal model for arbitrary pseudoeffecitve $\mathbb{R}$-divisor $-(K_{X}+\Delta)$.
Assume that $X$ is $\mathbb{Q}$-factorial, and let $-(K_{X}+\Delta)=P+N$ be the divisorial Zariski decomposition. Take a log resolution $f:Y\to X$ of $(X,\Delta+N)$. Then $E=N_{\sigma}(-f^{\ast}(K_{X}+\Delta))-f^{\ast}N$ is an $f$-exceptional divisor. By Lemma 3.3, we have $E\geq 0$. That is, $f^{\ast}N\leq N_{\sigma}(-f^{\ast}(K_{X}+\Delta))$ and by Lemma 3.5, we have
(4.6) $N_{\sigma}(f^{\ast}P)=N_{\sigma}(-f^{\ast}(K_{X}+\Delta)-f^{\ast}N)=N_{\sigma}(-f^{\ast}(K_{X}+\Delta))-f^{\ast}N=E.$
Write
$K_{Y}+\Delta_{Y}=f^{\ast}(K_{X}+\Delta).$
Then by the construction of $E$, we have
(4.7) $K_{Y}+\Delta_{Y}+N_{\sigma}(-f^{\ast}(K_{X}+\Delta))=f^{\ast}(K_{X}+\Delta+N)+E.$
Note that $-(K_{X}+\Delta+N)=P$ is pseudoeffective. Furthermore, for any prime divisor $F$ on $Y$, by (4.6) and (4.7), we have
$\bar{a}(F;X,\Delta+N)$ $=1-\operatorname{mult}_{F}(\Delta_{Y}+N_{\sigma}(-f^{\ast}(K_{X}+\Delta))-E)-\operatorname{mult}_{F}E$
$=1-\operatorname{mult}_{F}(\Delta_{Y}+N_{\sigma}(-f^{\ast}(K_{X}+\Delta)))$
$=\bar{a}(F;X,\Delta).$
It follows that $(X,\Delta)$ is pklt if and only if $(X,\Delta+N)$ is pklt.
Now, suppose that $(X,\Delta+N)$ is pklt. Let $\phi:X\dashrightarrow X^{\prime}$ be the birational contraction such that $(X^{\prime},\Delta^{\prime}+N^{\prime})$ is a $-(K_{X}+\Delta+N)$-minimal model and let $f:Y\to X$ and $g:Y\to X^{\prime}$ be a common resolution such that $g=\phi\circ f$. We may assume that $f$ is a log resolution of $(X,\Delta+N)$. Write
$-f^{\ast}(K_{X}+\Delta+N)=-g^{\ast}(K_{X^{\prime}}+\Delta^{\prime}+N^{\prime})+\Phi,$
where $\Phi$ is a $g$-exceptional divisor. As we have seen in Remark 4.2, $\Phi$ is an effective divisor and $N_{\sigma}(-f^{\ast}(K_{X}+\Delta+N))=\Phi$. On the other hand, by (4.6),
$N_{\sigma}(-f^{\ast}(K_{X}+\Delta+N))=N_{\sigma}(f^{\ast}P)=E.$
Hence $E=\Phi$. Note that
$P_{\sigma}(-f^{\ast}(K_{X}+\Delta))$ $=f^{\ast}P-E$
$=-f^{\ast}(K_{X}+\Delta+N)-\Phi$
$=-g^{\ast}(K_{X^{\prime}}+\Delta^{\prime}+N^{\prime}).$
Thus the positive part $P_{\sigma}(-f^{\ast}(K_{X}+\Delta))$ is nef, that is, $-(K_{X}+\Delta)$ admits a birational Zariski decomposition. Now, by Theorem 1.3, we can conclude that $(X,\Delta)$ has a $-(K_{X}+\Delta)$-minimal model.